# Geometry of efficient weight vectors

### v1

@@ 3 Results
$\mathbf{A_{21}}=\begin{pmatrix}1&1/a&d&e\\ a&1&b&c\\ 1/d&1/b&1&f\\ 1/e&1/c&1/f&1\end{pmatrix}$
Let us assume that in $\mathbf{A}$
  • for triad (1,2,3) $ad>b$,
  • for triad (2,3,4) $df>e$,
  • for triad (1,3,4) $bf>c$,
  • for triad (1,2,4) $ae>c$.
Then, in $\mathbf{A_{21}}$
  • for triad (1,2,3) $(1/a)\cdot b<d$,
  • for triad (2,3,4) $bf>c$,
  • for triad (1,3,4) $df>e$,
  • for triad (1,2,4) $(1/a)\cdot c<e$.
@@
Any $4\times 4$ pairwise comparison matrix can be transformed with appropriate reindexing and multiplication of weights of alternatives to one of the following forms:
$\begin{pmatrix}1&1&b^{\prime\prime}&c^{\prime\prime}\\ 1&1&1&e^{\prime\prime}\\ 1/b^{\prime\prime}&1&1&1\\ 1/c^{\prime\prime}&1/e^{\prime\prime}&1&1\end{pmatrix},$
Case 1: $1\cdot 1<b^{\prime\prime}$, $1\cdot 1<e^{\prime\prime}$, $b^{\prime\prime}\cdot 1<c^{\prime\prime}$, $1\cdot e^{\prime\prime}<c^{\prime\prime}$.
Case 2: $1\cdot 1<b^{\prime\prime}$, $1\cdot 1<e^{\prime\prime}$, $b^{\prime\prime}\cdot 1<c^{\prime\prime}$, $1\cdot e^{\prime\prime}>c^{\prime\prime}$.
Proof.
Apply Proposition 2 to the pairwise comparison matrix in Proposition 1. ∎
Corollary 3 means that it is enough to examine these two cases, and show the geometric nature of the set of efficient weight vectors, and the directions of the arcs in the strongly connected BCC graphs.
First, let us consider Case 1. Szádoczki and Bozóki, (2024) showed that the set of efficient weight vectors for $4$ alternatives geometrically consists of the union of three tetrahedra (denoted by $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) associated with the labeled Hamiltonian cycles of BCC graphs presented in Figure 2, where for a given PCM always only one of the two directions is possible for each cycle.
[…]
Figure 2: The labeled Hamiltonian cycles of BCC graphs associated with tetrahedra $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$. Each tetrahedron is related to the cycles below them, from which always only one is possible for a given PCM.
As shown in Figure 1 there is a cutting plane for every element of the pairwise comparison matrix (aside from the elements coming from the reciprocity property). However, each of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$ is determined by exactly four planes. Figures 3, 4, and 5 show $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$, respectively, with the planes determining them, while Figure 6 present the set of efficient weight vectors (the union of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) for the following illustrative example matrix ($b^{\prime\prime}=2$, $c^{\prime\prime}=4$, $e^{\prime\prime}=3$):
$\mathbf{A^{\prime\prime}}=\begin{pmatrix}1&1&2&4\\ 1&1&1&3\\ 1/2&1&1&1\\ 1/4&1/3&1&1\end{pmatrix}.$
Note that in Figures 3, and 4 plane $h_{23}$ (denoted by dark green color) is barely visible from this angle, as it connects $V_{1}$, $V_{4}$ and a point of the $V_{2}V_{3}$ edge that happens to coincide with a point from the $V_{1}V_{4}$ edge, thus from this plane only the $V_{1}V_{4}$ edge can be observed. This is true for all the latter figures containing $h_{23}$ as well.
@@
Figure 3: The first tetrahedron of efficient weight vectors ($\mathcal{T}_{1}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{14},h_{23}$, and $h_{34}$) for the Case 1 example.
@@
Figure 4: The second tetrahedron of efficient weight vectors ($\mathcal{T}_{2}$), the associated Hamiltonian cycle, and the determining planes ($h_{13},h_{14},h_{23}$, and $h_{24}$) for the Case 1 example.
@@
Figure 5: The third tetrahedron of efficient weight vectors ($\mathcal{T}_{3}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{13},h_{24}$, and $h_{34}$) for the Case 1 example.

### v2

@@ 3 Results
Case 1A:  $a_{12}a_{23}a_{34}a_{41}<1$,  $a_{14}a_{42}a_{23}a_{31}<1$,  $a_{13}a_{34}a_{42}a_{21}<1$.
This case needs no rearrangement, the original $(1,2,3,4)$ satisfies the criteria. Note, that $(1,3,4,2)$ and $(1,4,2,3)$ will also satisfy the criteria in this case, since
$\mathbf{A}_{1342}=\begin{pmatrix}1&a_{13}&a_{14}&a_{12}\\ 1/a_{13}&1&a_{34}&1/a_{23}\\ 1/a_{14}&1/a_{34}&1&1/a_{24}\\ 1/a_{12}&a_{23}&a_{24}&1\end{pmatrix},\ \mathbf{A}_{1423}=\begin{pmatrix}1&a_{14}&a_{12}&a_{13}\\ 1/a_{14}&1&1/a_{24}&1/a_{34}\\ 1/a_{12}&a_{24}&1&a_{23}\\ 1/a_{13}&a_{34}&1/a_{23}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$a_{13}\cdot a_{34}\cdot 1/a_{24}\cdot 1/a_{12}<1,\quad a_{12}\cdot a_{23}\cdot a_{34}\cdot 1/a_{14}<1,\quad a_{14}\cdot 1/a_{24}\cdot a_{23}\cdot 1/a_{13}<1$
for matrix $\mathbf{A}_{1342}$ and they are
$a_{14}\cdot 1/a_{24}\cdot a_{23}\cdot 1/a_{13}<1,\quad a_{13}\cdot a_{34}\cdot 1/a_{24}\cdot 1/a_{12}<1,\quad a_{12}\cdot a_{23}\cdot a_{34}\cdot 1/a_{14}<1$
for matrix $\mathbf{A}_{1423}$.
The arrangements for this case are when alternative $1$ is put as first, and the directed triad $(2,3,4)$ after that, at any starting point in the latter.
Case 1B:  $a_{12}a_{23}a_{34}a_{41}>1$,  $a_{14}a_{42}a_{23}a_{31}>1$,  $a_{13}a_{34}a_{42}a_{21}>1$.
Rearrangements $(1,2,4,3)$, $(1,3,2,4)$, and $(1,4,3,2)$ satisfy the requirement.
$\mathbf{A}_{1243}=\begin{pmatrix}1&a_{12}&a_{14}&a_{13}\\ 1/a_{12}&1&a_{24}&a_{23}\\ 1/a_{14}&1/a_{24}&1&1/a_{34}\\ 1/a_{13}&1/a_{23}&a_{34}&1\end{pmatrix},\ \mathbf{A}_{1324}=\begin{pmatrix}1&a_{13}&a_{12}&a_{14}\\ 1/a_{13}&1&1/a_{23}&a_{34}\\ 1/a_{12}&a_{23}&1&a_{24}\\ 1/a_{14}&1/a_{34}&1/a_{24}&1\\ \end{pmatrix},$
$\mathbf{A}_{1432}=\begin{pmatrix}1&a_{14}&a_{13}&a_{12}\\ 1/a_{14}&1&1/a_{34}&1/a_{24}\\ 1/a_{13}&a_{34}&1&1/a_{23}\\ 1/a_{12}&a_{24}&a_{23}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$ , and $(1,3,4,2,1)$ are
$a_{12}\cdot a_{24}\cdot 1/a_{34}\cdot 1/a_{13}<1,\quad a_{13}\cdot 1/a_{23}\cdot a_{24}\cdot 1/a_{14}<1,\quad a_{14}\cdot 1/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}<1$
for matrix $\mathbf{A}_{1243}$, while they are
$a_{13}\cdot 1/a_{23}\cdot a_{24}\cdot 1/a_{14}<1,\quad a_{14}\cdot 1/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}<1,\quad a_{12}\cdot a_{24}\cdot 1/a_{34}\cdot 1/a_{13}<1$
for matrix $\mathbf{A}_{1324}$ and they are
$a_{14}\cdot 1/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}<1,\quad a_{12}\cdot a_{24}\cdot 1/a_{34}\cdot 1/a_{13}<1,\quad a_{13}\cdot 1/a_{23}\cdot a_{24}\cdot 1/a_{14}<1$
for matrix $\mathbf{A}_{1432}$.
The arrangements for this case are when alternative $1$ is put as first and the directed triad $(4,3,2)$ after that, at any starting point in the latter.
Case 2A:  $a_{12}a_{23}a_{34}a_{41}<1$,  $a_{14}a_{42}a_{23}a_{31}>1$,  $a_{13}a_{34}a_{42}a_{21}<1$.
Rearrangements $(2,1,3,4)$, $(2,3,4,1)$, and $(2,4,1,3)$ satisfy the requirement.
$\mathbf{A}_{2134}=\begin{pmatrix}1&1/a_{12}&a_{23}&a_{24}\\ a_{12}&1&a_{13}&a_{14}\\ 1/a_{23}&1/a_{13}&1&a_{34}\\ 1/a_{24}&1/a_{14}&1/a_{34}&1\\ \end{pmatrix},\ \mathbf{A}_{2341}=\begin{pmatrix}1&a_{23}&a_{24}&1/a_{12}\\ 1/a_{23}&1&a_{34}&1/a_{13}\\ 1/a_{24}&1/a_{34}&1&1/a_{14}\\ a_{12}&a_{13}&a_{14}&1\\ \end{pmatrix},$
$\mathbf{A}_{2413}=\begin{pmatrix}1&a_{24}&1/a_{12}&a_{23}\\ 1/a_{24}&1&1/a_{14}&1/a_{34}\\ a_{12}&a_{14}&1&a_{13}\\ 1/a_{23}&a_{34}&1/a_{13}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{12}\cdot a_{13}\cdot a_{34}\cdot 1/a_{24}<1,\quad a_{24}\cdot 1/a_{14}\cdot a_{13}\cdot 1/a_{23}<1,\quad a_{23}\cdot a_{34}\cdot 1/a_{14}\cdot a_{12}<$ 1
for matrix $\mathbf{A}_{2134}$, while they are
$a_{23}\cdot a_{34}\cdot 1/a_{14}\cdot a_{12}<1,\quad 1/a_{12}\cdot a_{13}\cdot a_{34}\cdot 1/a_{24}<1,\quad a_{24}\cdot 1/a_{14}\cdot a_{13}\cdot 1/a_{23}<1$
for matrix $\mathbf{A}_{2341}$ and they are
$a_{24}\cdot 1/a_{14}\cdot a_{13}\cdot 1/a_{23}<1,\quad a_{23}\cdot a_{34}\cdot 1/a_{14}\cdot a_{12}<1,\quad 1/a_{12}\cdot a_{13}\cdot a_{34}\cdot 1/a_{24}<1$
for matrix $\mathbf{A}_{2413}$.
The arrangements for this case are when alternative $2$ is put as first and the directed triad $(1,3,4)$ after that, at any starting point in the latter.
Case 2B:  $a_{12}a_{23}a_{34}a_{41}>1$,  $a_{14}a_{42}a_{23}a_{31}<1$,  $a_{13}a_{34}a_{42}a_{21}>1$.
Rearrangements $(2,1,4,3)$, $(2,4,3,1)$, and $(2,3,1,4)$ satisfy the requirement.
$\mathbf{A}_{2143}=\begin{pmatrix}1&1/a_{12}&a_{24}&a_{23}\\ a_{12}&1&a_{14}&a_{13}\\ 1/a_{24}&1/a_{14}&1&1/a_{34}\\ 1/a_{23}&1/a_{13}&a_{34}&1\\ \end{pmatrix},\ \mathbf{A}_{2431}=\begin{pmatrix}1&a_{24}&a_{23}&1/a_{12}\\ 1/a_{24}&1&1/a_{34}&1/a_{14}\\ 1/a_{23}&a_{34}&1&1/a_{13}\\ a_{12}&a_{14}&a_{13}&1\\ \end{pmatrix},$
$\mathbf{A}_{2314}=\begin{pmatrix}1&a_{23}&1/a_{12}&a_{24}\\ 1/a_{23}&1&1/a_{13}&a_{34}\\ a_{12}&a_{13}&1&a_{14}\\ 1/a_{24}&1/a_{34}&1/a_{14}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{12}\cdot a_{14}\cdot 1/a_{34}\cdot 1/a_{23}<1,\quad a_{23}\cdot 1/a_{13}\cdot a_{14}\cdot 1/a_{24}<1,\quad a_{24}\cdot 1/a_{34}\cdot 1/a_{13}\cdot a_{12}<$ 1
for matrix $\mathbf{A}_{2143}$, while they are
$a_{24}\cdot 1/a_{34}\cdot 1/a_{13}\cdot a_{12}<1,\quad 1/a_{12}\cdot a_{14}\cdot 1/a_{34}\cdot 1/a_{23}<1,\quad a_{23}\cdot 1/a_{13}\cdot a_{14}\cdot 1/a_{24}<1$
for matrix $\mathbf{A}_{2431}$ and they are
$a_{23}\cdot 1/a_{13}\cdot a_{14}\cdot 1/a_{24}<1,\quad a_{24}\cdot 1/a_{34}\cdot 1/a_{13}\cdot a_{12}<1,\quad 1/a_{12}\cdot a_{14}\cdot 1/a_{34}\cdot 1/a_{23}<1$
for matrix $\mathbf{A}_{2314}$.
The arrangements for this case are when alternative $2$ is put as first and the directed triad $(4,3,1)$ after that, at any starting point in the latter.
Case 3A:  $a_{12}a_{23}a_{34}a_{41}>1$,  $a_{14}a_{42}a_{23}a_{31}<1$,  $a_{13}a_{34}a_{42}a_{21}<1$.
Rearrangements $(3,1,4,2)$, $(3,4,2,1)$, and $(3,2,1,4)$ satisfy the requirement.
$\mathbf{A}_{3142}=\begin{pmatrix}1&1/a_{13}&a_{34}&1/a_{23}\\ a_{13}&1&a_{14}&a_{12}\\ 1/a_{34}&1/a_{14}&1&1/a_{24}\\ a_{23}&1/a_{12}&a_{24}&1\\ \end{pmatrix},\ \mathbf{A}_{3421}=\begin{pmatrix}1&a_{34}&1/a_{23}&1/a_{13}\\ 1/a_{34}&1&1/a_{24}&1/a_{14}\\ a_{23}&a_{24}&1&1/a_{12}\\ a_{13}&a_{14}&a_{12}&1\\ \end{pmatrix},$
$\mathbf{A}_{3214}=\begin{pmatrix}1&1/a_{23}&1/a_{13}&a_{34}\\ a_{23}&1&1/a_{12}&a_{24}\\ a_{13}&a_{12}&1&a_{14}\\ 1/a_{34}&1/a_{24}&1/a_{14}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{13}\cdot a_{14}\cdot 1/a_{24}\cdot a_{23}<1,\quad 1/a_{23}\cdot 1/a_{12}\cdot a_{14}\cdot 1/a_{34}<1,\quad a_{34}\cdot 1/a_{24}\cdot 1/a_{12}\cdot a_{13}<$ 1
for matrix $\mathbf{A}_{3142}$, while they are
$a_{34}\cdot 1/a_{24}\cdot 1/a_{12}\cdot a_{13}<1,\quad 1/a_{13}\cdot a_{14}\cdot 1/a_{24}\cdot a_{23}<1,\quad 1/a_{23}\cdot 1/a_{12}\cdot a_{14}\cdot 1/a_{34}<1$
for matrix $\mathbf{A}_{3421}$ and they are
$11/a_{23}\cdot 1/a_{12}\cdot a_{14}\cdot 1/a_{34}<1,\quad a_{34}\cdot 1/a_{24}\cdot 1/a_{12}\cdot a_{13}<1,\quad 1/a_{13}\cdot a_{14}\cdot 1/a_{24}\cdot a_{23}<$ 1
for matrix $\mathbf{A}_{3214}$.
The arrangements for this case are when alternative $3$ is put as first and the directed triad $(4,2,1)$ after that, at any starting point in the latter.
Case 3B:  $a_{12}a_{23}a_{34}a_{41}<1$,  $a_{14}a_{42}a_{23}a_{31}>1$,  $a_{13}a_{34}a_{42}a_{21}>1$.
Rearrangements $(3,1,2,4)$, $(3,2,4,1)$, and $(3,4,1,2)$ satisfy the requirement.
$\mathbf{A}_{3124}=\begin{pmatrix}1&1/a_{13}&1/a_{23}&a_{34}\\ a_{13}&1&a_{12}&a_{14}\\ a_{23}&1/a_{12}&1&a_{24}\\ 1/a_{34}&1/a_{14}&1/a_{24}&1\\ \end{pmatrix},\ \mathbf{A}_{3241}=\begin{pmatrix}1&1/a_{23}&a_{34}&1/a_{13}\\ a_{23}&1&a_{24}&1/a_{12}\\ 1/a_{34}&1/a_{24}&1&1/a_{14}\\ a_{13}&a_{12}&a_{14}&1\\ \end{pmatrix},$
$\mathbf{A}_{3412}=\begin{pmatrix}1&a_{34}&1/a_{13}&1/a_{23}\\ 1/a_{34}&1&1/a_{14}&1/a_{24}\\ a_{13}&a_{14}&1&a_{12}\\ a_{23}&a_{24}&1/a_{12}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{13}\cdot a_{12}\cdot a_{24}\cdot 1/a_{34}<1,\quad a_{34}\cdot 1/a_{14}\cdot a_{12}\cdot a_{23}<1,\quad 1/a_{23}\cdot a_{24}\cdot 1/a_{14}\cdot a_{13}<$ 1
for matrix $\mathbf{A}_{3124}$, while they are
$11/a_{23}\cdot a_{24}\cdot 1/a_{14}\cdot a_{13}<1,\quad 1/a_{13}\cdot a_{12}\cdot a_{24}\cdot 1/a_{34}<1,\quad a_{34}\cdot 1/a_{14}\cdot a_{12}\cdot a_{23}<$ 1
for matrix $\mathbf{A}_{3241}$ and they are
$a_{34}\cdot 1/a_{14}\cdot a_{12}\cdot a_{23}<1,\quad 1/a_{23}\cdot a_{24}\cdot 1/a_{14}\cdot a_{13}<1,\quad 1/a_{13}\cdot a_{12}\cdot a_{24}\cdot 1/a_{34}<1$
for matrix $\mathbf{A}_{3412}$.
The arrangements for this case are when alternative $3$ is put as first and the directed triad $(1,2,4)$ after that, at any starting point in the latter.
Case 4A:  $a_{12}a_{23}a_{34}a_{41}<1$,  $a_{14}a_{42}a_{23}a_{31}<1$,  $a_{13}a_{34}a_{42}a_{21}>1$.
Rearrangements $(4,1,2,3)$, $(4,2,3,1)$, and $(4,3,1,2)$ satisfy the requirement.
$\mathbf{A}_{4123}=\begin{pmatrix}1&1/a_{14}&1/a_{24}&1/a_{34}\\ a_{14}&1&a_{12}&a_{13}\\ a_{24}&1/a_{12}&1&a_{23}\\ a_{34}&1/a_{13}&1/a_{23}&1\\ \end{pmatrix},\ \mathbf{A}_{4231}=\begin{pmatrix}1&1/a_{24}&1/a_{34}&1/a_{14}\\ a_{24}&1&a_{23}&1/a_{12}\\ a_{34}&1/a_{23}&1&1/a_{13}\\ a_{14}&a_{12}&a_{13}&1\\ \end{pmatrix},$
$\mathbf{A}_{4312}=\begin{pmatrix}1&1/a_{34}&1/a_{14}&1/a_{24}\\ a_{34}&1&1/a_{13}&1/a_{23}\\ a_{14}&a_{13}&1&a_{12}\\ a_{24}&a_{23}&1/a_{12}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{14}\cdot a_{12}\cdot a_{23}\cdot a_{34}<1,\quad 1/a_{34}\cdot 1/a_{13}\cdot a_{12}\cdot a_{24}<1,\quad 1/a_{24}\cdot a_{23}\cdot 1/a_{13}\cdot a_{14}<$ 1
for matrix $\mathbf{A}_{4123}$, while they are
$11/a_{24}\cdot a_{23}\cdot 1/a_{13}\cdot a_{14}<1,\quad 1/a_{14}\cdot a_{12}\cdot a_{23}\cdot a_{34}<1,\quad 1/a_{34}\cdot 1/a_{13}\cdot a_{12}\cdot a_{24}<$ 1
for matrix $\mathbf{A}_{4231}$ and they are
$11/a_{34}\cdot 1/a_{13}\cdot a_{12}\cdot a_{24}<1,\quad 1/a_{24}\cdot a_{23}\cdot 1/a_{13}\cdot a_{14}<1,\quad 1/a_{14}\cdot a_{12}\cdot a_{23}\cdot a_{34}<$ 1
for matrix $\mathbf{A}_{4312}$.
The arrangements for this case are when alternative $4$ is put as first and the directed triad $(1,2,3)$ after that, at any starting point in the latter.
Case 4B:  $a_{12}a_{23}a_{34}a_{41}>1$,  $a_{14}a_{42}a_{23}a_{31}>1$,  $a_{13}a_{34}a_{42}a_{21}<1$.
Rearrangements $(4,1,3,2)$, $(4,2,1,3)$, and $(4,3,2,1)$ satisfy the requirement.
$\mathbf{A}_{4132}=\begin{pmatrix}1&1/a_{14}&1/a_{34}&1/a_{24}\\ a_{14}&1&a_{13}&a_{12}\\ a_{34}&1/a_{13}&1&1/a_{23}\\ a_{24}&1/a_{12}&a_{23}&1\\ \end{pmatrix},\ \mathbf{A}_{4213}=\begin{pmatrix}1&1/a_{24}&1/a_{14}&1/a_{34}\\ a_{24}&1&1/a_{12}&a_{23}\\ a_{14}&a_{12}&1&a_{13}\\ a_{34}&1/a_{23}&1/a_{13}&1\\ \end{pmatrix},$
$\mathbf{A}_{4321}=\begin{pmatrix}1&1/a_{34}&1/a_{24}&1/a_{14}\\ a_{34}&1&1/a_{23}&1/a_{13}\\ a_{24}&a_{23}&1&1/a_{12}\\ a_{14}&a_{13}&a_{12}&1\\ \end{pmatrix},$
and the product of the elements along cycles $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$ are
$11/a_{14}\cdot a_{13}\cdot 1/a_{23}\cdot a_{24}<1,\quad 1/a_{24}\cdot 1/a_{12}\cdot a_{13}\cdot a_{34}<1,\quad 1/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}\cdot a_{14}<$ 1
for matrix $\mathbf{A}_{4132}$, while they are
$11/a_{24}\cdot 1/a_{12}\cdot a_{13}\cdot a_{34}<1,\quad 1/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}\cdot a_{14}<1,\quad 1/a_{14}\cdot a_{13}\cdot 1/a_{23}\cdot a_{24}<$ 1
for matrix $\mathbf{A}_{4213}$ and they are
$11/a_{34}\cdot 1/a_{23}\cdot 1/a_{12}\cdot a_{14}<1,\quad 1/a_{14}\cdot a_{13}\cdot 1/a_{23}\cdot a_{24}<1,\quad 1/a_{24}\cdot 1/a_{12}\cdot a_{13}\cdot a_{34}<$ 1
for matrix $\mathbf{A}_{4321}$.
The arrangements for this case are when alternative $4$ is put as first and the directed triad $(3,2,1)$ after that, at any starting point in the latter. ∎
That means any PCM can be rearranged in a way that the directed $4$-cycles have a given orientation. In order to prove that the set of efficient weight vectors is determined by three tetrahedra ($\mathcal{T}_{1}$, $\mathcal{T}_{2}$, and $\mathcal{T}_{3}$) a number of lemmas will be necessary.
Lemma 1.
For a $4\times 4$ pairwise comparison matrix $\mathbf{A}$ the weight vectors calculated from any of its incomplete submatrices which can be represented by a spanning tree will be efficient, and each will estimate three elements of $\mathbf{A}$ perfectly.
Proof.
An incomplete pairwise comparison submatrix of $\mathbf{A}$ that can be represented by a spanning tree contains three comparisons and there are no $k$-cycles ($k\geq 3$) in it. That means there cannot be inconsistency between them, the given submatrix can be complemented in a consistent way. Hence the $\mathbf{w}$ weight vector calculated from it will estimate the three given elements perfectly.
There cannot be a $\mathbf{w}^{\prime}$ that estimates these three elements perfectly and at least one other element better than $\mathbf{w}$, as the single normalized weight vector that can estimate three elements of the matrix (and thus the consistent PCM calculated from them) perfectly is $\mathbf{w}$. Thus, $\mathbf{w}$ is efficient for $\mathbf{A}$. ∎
Lemma 2.
For the BCC graphs of a $4\times 4$ pairwise comparison matrix $\mathbf{A}$ (that satisfies Assumption 1), from the six labeled Hamiltonian directed cycles exactly three are possible, which are the three labeled Hamiltonian undirected cycles with exactly one orientation each.
Proof.
There are three labeled Hamiltonian undirected cycles for four alternatives: $(1,2,3,4,1)$, $(1,4,2,3,1)$, and $(1,3,4,2,1)$. Let us consider the $(1,2,3,4,1)$ cycle, which has two possible orientations.
Without the loss of generality, let us assume that there exists a weight vector $\mathbf{w}$ for which the orientation of the $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ BCC graph is $(1,2,3,4,1)$. That means the following:
$w_{1}/w_{2}\geq a_{12}\iff w_{1}\geq a_{12}w_{2}$
$w_{2}/w_{3}\geq a_{23}\iff w_{2}\geq a_{23}w_{3}$
$w_{3}/w_{4}\geq a_{34}\iff w_{3}\geq a_{34}w_{4}$
$w_{4}/w_{1}\geq a_{41}\iff w_{4}\geq a_{41}w_{1}$
Writing Inequality 3 into 3, and then 3 into the result, and finally 3 into the result, we get:
$w_{1}\geq a_{12}a_{23}a_{34}a_{41}w_{1}\iff 1\geq a_{12}a_{23}a_{34}a_{41}$
If we assume a weight vector with the other orientation of the BCC graph, then Inequality 3 will be reversed. The only possible way to get BCC graphs with both orientations is to get an equality in 3. However, that means the $(1,2,3,4,1)$ $4$-cycle is consistent, and that violates Assumption 1.
The same logic can be used for the other two $4$-cycles. ∎
Proposition 2.
For any $4\times 4$ pairwise comparison matrix $\mathbf{A}$ (that satisfies Assumption 1), there exists $\mathbf{A^{\prime\prime}}$ obtained from $\mathbf{A}$ by appropriate rearrangements, for which the set of efficient weight vectors can be constructed as the union of three tetrahedra, where each tetrahedron is determined by four weight vectors, each calculated from an incomplete pairwise comparison submatrix of $\mathbf{A^{\prime\prime}}$ with a path-type spanning tree representing graph. Additionally, for each efficient weight vector one of the following directed cycles is present in its BCC graph: (1,2,3,4,1), (1,4,2,3,1), or (1,3,4,2,1).
Proof.
Based on Corollary 1 the BCC graph of any efficient weight vector must contain a directed Hamiltonian cycle.
According to Lemma 2 for a given $4\times 4$ PCM $\mathbf{A}$, exactly three directed Hamiltonian cycles are possible.
Based on Lemma 1 the weight vectors calculated from an incomplete submatrix of $\mathbf{A}$ with a spanning tree representing graph must be efficient, and
  • each estimates three elements of $\mathbf{A}$ perfectly
[…]
  • they are at the intersection of three of the possible six cutting planes $h_{12},\ldots,h_{34}$
[…]
  • in their BCC graphs there are arcs in both directions between three pairs of vertices.
Apply the reindexing of Proposition 1 to $\mathbf{A}$ to gain $\mathbf{A}^{\prime\prime}$, where the three possible directed Hamiltonian cycles’ orientations are: $(1,2,3,4,1)$, $(1,4,2,3,1)$ and $(1,3,4,2,1)$.
Let us consider the $(1,2,3,4,1)$ cycle. There are four incomplete pairwise comparison submatrix of $\mathbf{A}^{\prime\prime}$ with a path spanning tree representing graph that includes only comparisons that are related to this cycle (denoted by $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$ respectively):
  • $T_{1}:(1,2,3,4)$, excluding the comparison of $(1,4)$,
  • $T_{2}:(2,3,4,1)$, excluding the comparison of $(1,2)$,
  • $T_{3}:(3,4,1,2)$, excluding the comparison of $(2,3)$,
  • $T_{4}:(4,1,2,3)$, excluding the comparison of $(3,4)$.
Based on Assumption 1 there are no consistent $k$-cycles ($k\geq 3$) in $\mathbf{A}^{\prime\prime}$, thus the weight vectors calculated from $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$ do not correspond with each other.
These are efficient, and as four different points, they determine a tetrahedron in the space of normalized ($\sum_{i=1}^{4}w_{i}=1$) $4$-dimensional weight vectors.
The edges of this tetrahedron also only contain efficient weight vectors. Without the loss of generality, let us consider edge $T_{1}T_{2}$. It ensures that two comparisons $(2,3)$ and $(3,4)$ are estimated perfectly, and at each of $T_{1}$ and $T_{2}$ another third element is approximated accurately ($(1,2)$ and $(4,1)$ respectively). This means that on the $T_{1}T_{2}$ edge one of the $(1,2)$ and $(4,1)$ comparisons can only be estimated better with the worse estimation of the other one, thus all the weight vectors here are efficient. Similar argument can be made for all edges of the tetrahedron.
@@
For any $4\times 4$ pairwise comparison matrix $\mathbf{A}$ (that satisfies Assumption 1), every efficient weight vector can be written as the convex combination of the $4$ weight vectors determined by the path-type spanning tree representing graphs for one of the three cycles of (1,2,3,4,1), (1,4,2,3,1), or (1,3,4,2,1).
Remark 4.
The results of Proposition 1, Proposition 2, and Corollary 2 hold without Assumption 1 with appropriate modifications (allowing equaility as well). Thus, the main results are the same even if there are consistent cycles in the PCM, however, each consistent subset will only result in one possible efficient weight vector. For all the special cases with examples, see Appendix Appendix B.
Let us consider an example geometrically.
Example 3.
As mentioned before, each of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$ is determined by exactly four cutting planes. Figures 5, 6, and 7 show $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$, respectively, with the planes determining them, while Figure 8 presents the set of efficient weight vectors (the union of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) for the following illustrative example matrix ($a_{12}^{\prime\prime}=1$, $a_{13}^{\prime\prime}=5$, $a_{14}^{\prime\prime}=7$, $a_{23}^{\prime\prime}=2$, $a_{24}^{\prime\prime}=8$, and $a_{34}^{\prime\prime}=1/3$):
$\mathbf{A^{\prime\prime}}=\begin{pmatrix}1&1&5&7\\ 1&1&2&8\\ 1/5&1/2&1&1/3\\ 1/7&1/8&3&1\end{pmatrix}.$
@@
Figure 5: The first tetrahedron of efficient weight vectors ($\mathcal{T}_{1}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{14},h_{23}$, and $h_{34}$) for the example.
@@
Figure 6: The second tetrahedron of efficient weight vectors ($\mathcal{T}_{2}$), the associated Hamiltonian cycle, and the determining planes ($h_{13},h_{14},h_{23}$, and $h_{24}$) for the example.
@@
Figure 7: The third tetrahedron of efficient weight vectors ($\mathcal{T}_{3}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{13},h_{24}$, and $h_{34}$) for the example.